OpenStudy (shayaan_mustafa):

f(x)=2*sqrt(x-1) and g(x)=sqrt(x-1) compute the functions and find domains of the following: 1) f(x)+g(x) 2) f(x)-g(x) 3) f(x)*g(x) 4) f(x)/g(x)

OpenStudy (anonymous):

Replace sqrt(x-1) with "y" and everything is trivial, except one that results in y squared and will need to be evaluated. Assuming these are real functions, where is the sqrt function defined?

myininaya (myininaya):

\[f(x)+g(x)=2 \sqrt{x-1}+\sqrt{x-1} \] f+g will be continuous when f and g are continuous x-1 has to be positive or zero, right?

OpenStudy (shayaan_mustafa):

yes right myininaya. f(x)+g(x)=3sqrt(x-1) (+inf,1)right?

myininaya (myininaya):

\[x-1 \ge 0\] \[ x \ge 1 \] Which it around (1,inf) since inf>1

myininaya (myininaya):

And also include 1 so we say [1,inf)

OpenStudy (shayaan_mustafa):

no no wait wait. [1,inf) means 1 is included and inf is not. but why don't inf?

myininaya (myininaya):

because inf is not actually a number

myininaya (myininaya):

also \[\sqrt{1-1}=\sqrt{0}=0 \text{; 0 exists!}\]

myininaya (myininaya):

so we include 1

OpenStudy (shayaan_mustafa):

oh I see. so thus inf can never be a domain of any kind of function. right??

myininaya (myininaya):

Right when we say [1,inf) this just means we want to include 1 and every single number greater than 1 We never include infinity because it is not a number

OpenStudy (shayaan_mustafa):

ok ok. fine.. 2)sqrt(x-1) domain:[1,inf) 3)2x-2 what would be this domain?

OpenStudy (shayaan_mustafa):

OK. This makes sense.

myininaya (myininaya):

\[f(x) \cdot g(x)=2 \sqrt{x-1} \cdot \sqrt{x-1}\] Domain here is still the same We want to \[\sqrt{x-1} \text{ to be a real number}\] In order to have that that is a real number we must again have that \[x-1 \ge 0\] Which means \[x \ge 1 \] Which in interval notation that means [1,inf)

OpenStudy (shayaan_mustafa):

no no. look. why domain remain same?? this function is the multiple of two different function. so its domain different other than f(x)+g(x) and f(x)-g(x). explain.

myininaya (myininaya):

We want both f and g to be real right?

OpenStudy (shayaan_mustafa):

ok. so we look individually though we are computing any function?

OpenStudy (shayaan_mustafa):

4)2 domain plz...

myininaya (myininaya):

ok so 4 4) this time we have the \[\sqrt{x-1}\] on the bottom

OpenStudy (shayaan_mustafa):

yes.

myininaya (myininaya):

\[\frac{f(x)}{g(x)}=\frac{2 \sqrt{x-1}}{\sqrt{x-1}}\] We don't want the bottom to be zero

myininaya (myininaya):

so we don't want to include 1

myininaya (myininaya):

but we also don't want x-1<0 we want x-1>0 which means x>1 which means in interval notation we have (1,inf)

OpenStudy (shayaan_mustafa):

so it can never be 1 and any -ve value. right?

OpenStudy (shayaan_mustafa):

yes (1, inf)

myininaya (myininaya):

:)

OpenStudy (shayaan_mustafa):

ok w8 teacher. thnx..

OpenStudy (shayaan_mustafa):

this question was from Howard Anton calculus. But book's answer is different.

myininaya (myininaya):

awesome stuff :)

OpenStudy (shayaan_mustafa):

teacher. are you there?

myininaya (myininaya):

you have a question?

OpenStudy (shayaan_mustafa):

yes.